# Frozen Transformers in Language Models Are Effective Visual Encoder Layers

### v1

@@ B.2 Breakdown Metrics for 3D VQA
We additionally provide the full metrics on SQA3D in Table B, supplementing Table 4(b). As shown in the table, adding a frozen transformer improves the performance on the main metric and most of the analytical metrics.
@@
SQA3D + LLaMA 48.09 89.03 34.27 67.05 48.17 68.34 43.87 45.64
Table B: Performance of SQA3D dataset and adding language transformers on 3D Question Answering (QA). Adding LLaMA achieves best performance. EM@1 and EM@10 means Top-1 and Top-10 Exact Match (Accuracy) metric. “What”, “Is”, “How”, “Can”, “Which”, “Others” are detailed breakdown of question types reported in EM@1.
C Details of Implementations
@@ C.2 Point Cloud Classification
In this section, we describe the implementation details of the pointcloud classification method presented in Sec. 4.2. For optimization, we use AdamW (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) optimizer and a cosine annealing learning rate schedule (Loshchilov & Hutter, 2016). To map the dimension between the PointBERT and LLaMa transformer tokens, we add two linear layers with a learning rate of 5e-5. The PointBERT backbone has a learning rate of 5e-4, consistent with the original setting in (Yu et al., 2021).. We finetune our model for 300 epochs on both the ScanObjectNN (Uy et al., 2019) and ModelNet40 (Goyal et al., 2021) datasets. The training takes around 6-10 hours on 4$\times$×A200 GPUs
@@ C.3 Action Recognition
We investigate action recognition in Sec. 4.3 and provide more details of implementation here. Our setup strictly follows VideoMAE (Tong et al., 2022), where a ViT model is (1) pretrained by masked auto-encoding (He et al., 2022); then (2) finetuned for additional epochs. As stated in Sec. 4.3, we directly begin from the second step and inherit the parameters of self-attention blocks in ViT from pretrained VideoMAE models. During the training process, VideoMAE trains ViT-S for 40 epochs (5 epochs of warm-up) and ViT-B for 30 epochs (5 epochs of warm-up), where we adopt the same length of training. The optimizer is AdamW (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with the cosine annealing learning rate schedule (Loshchilov & Hutter, 2016).
In Sec. 4.3 and Table 3, our ViT-S and ViT-B performance is lower than the reported numbers in VideoMAE. This is because VideoMAE uses 32$\sim$64 GPUs during the finetuning stage and supports a much larger batch size compared to us, though we scale the learning rate according to the batch size as Goyal et al. (2017a). Concretely, VideoMAE adopts the batch size of 384 video clips, while our computational resource only supports a batch size of 24 clips and 12 clips for ViT-S and ViT-B, respectively. However, we control the setup between ViT and ViT-LLaMA for a fair comparison. Finally, the ViT-S-LLaMA and ViT-B-LLaMA experiments take around 3-4 days on 4$\times$×A100 GPUs.
@@ C.4 Motion Forecasting
Table C: Our implementation of VectorNet is better than their paper. Please note that this table uses the same single-modal setting (k=1), as VectorNet for comparison.
We evaluate the effects of frozen LLM transformers in Sec. 4.4 with motion forecasting. Since motion forecasting is a less well-known task in computer vision, we intuitively demonstrate its problem setting and modular architecture in Fig. G, where VectorNet (Gao et al., 2020) and mmTransformer (Liu et al., 2021) encode each lane or agent trajectory into a token embedding, then our LLM blocks process these tokens and feed them into the regression decoder.
Since VectorNet and mmTransformer have not released their training code, we reproduce their results on our own and achieve better or similar results as reported in the paper. As in Table C and Table D, the baseline used in our paper (Table 4) have comparable or even better performance compared to the original performance in the papers, which is critical for a fair and meaningful comparison.
@@
Table D: Our implementation of mmTransformer is comparable to the performance in their paper, with large advantages on the main metric of FDE.
During the training time, we separately train VectorNet or mmTransformer. VectorNet is a relatively simple architecture, so its training lasts 60 epochs, with cosine annealing learning rate schedule (Loshchilov & Hutter, 2016). We use the AdamW optimizer (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with a learning rate of 5e-4 and batch size equal to 32 samples. For mmTransformer, we train it with the same learning rate, batch size, and optimizer as VectorNet. The training lasts 32 epochs, where we drop the learning rate by 1/4 on epochs 20, 24, and 28. The training time for both models is around 2 days on one A100 GPU.
Figure G: Demonstration of typical motion forecasting design and our implementation. Motion forecasting models the trajectories of agents and lanes as polylines. Exiting models (Gao et al., 2020; Liu et al., 2021) using MLPs or transformers to convert the lanes and agent trajectories into token embeddings, then employ a decoder to regress the future trajectories. In our design, we treat either VectorNet (Gao et al., 2020) or mmTransformer (Liu et al., 2021) as general encoder, then insert the frozen LLM blocks to process their embeddings.
@@
This section provides more details on implementations and designs to supplement our discussion on 2D vision-language models in Sec. 4.5.1. Our experiments adopt the widely used METER (Dou et al., 2022) as the baseline and incorporate pretrained LLM transformers after its vision-language fusion module. In Fig. H, we intuitively illustrate the modular design of METER and the special place to insert our frozen LLM transformer and linear layers.
Conventionally, METER follows a two-stage training strategy: (1) pretrains the whole vision-language model (VLM) on a large combination of vision-language datasets, including COCO Lin et al. (2014), Conceptual Captions (Sharma et al., 2018), SBU Captions (Ordonez et al., 2011), and Visual Genome (Krishna et al., 2017); (2) finetuning on downstream tasks like visual question answering (VQA) or image-text retrieval. However, the first step of pretraining is computationally extensive, so we adopt the setup in Shi et al. (2023) by skipping the pretraining step and directly training on the target task. Specifically, we initialize the image encoder from CLIP-B/32 (Radford et al., 2021) and text encoder from RoBERTa (Liu et al., 2019), then finetune all the modules jointly expect for the LLM transformer $\mathbf{F}_{LM}$. Because of the initialization from CLIP and RoBERTa, our model is capable of predicting reasonably.
During the training stage, we strictly follow the same hyper-parameters and configurations on VQAv2 (Goyal et al., 2017c) and Flickr30k (Plummer et al., 2015) provided by METER. The most critical detail is that METER assigns different learning rates for each module. For example, the cross-modal fusion module and the decoder have larger learning rates compared to the pretrained image and text encoders. Similarly, our experiments set the learning rates of linear layers ($\mathbf{F}_{L}^{1}$ and $\mathbf{F}_{L}^{2}$) 10$\times$ the learning rate of the image encoder because they are randomly initialized. Finally, each training on VQAv2 and Flickr30k lasts for 10 epochs and around 1 day on 4$\times$×A100 GPUs.
Figure H: Intuitive illustration of METER (Dou et al., 2022) framework and how to insert the frozen language transformer (pink $\mathbf{F}_{LM}$) and linear layers (green $\mathbf{F}_{L}^{1}$ and $\mathbf{F}_{L}^{2}$) to process the visual tokens after vision-language fusion.
@@
This section provides more details of the dataset and training configurations of 3D vision language task. We conduct our experiments on SQA3D dataset (Ma et al., 2023), which contains $33.4k$ questions in $650$ unique ScanNet scenes. In addition to question answering (QA), the benchmark also requires the model to understand its situation (position, orientation, etc.) in the 3D scene as described by text. Hence it is called situated question answering (SQA). We use batch size 32 during our model training, and AdamW as our optimizer. The hidden size for each embedding token is 768. We train all parameters from scratch for 30 epochs, and decrease the learning rate by 10 times at 10, 15, and 20-th epoch. The model is trained on a single A100 GPU.
@@ C.7 Depths of LLM Layers
When varying the depth of transformer blocks in Sec. 5.2 and Fig. 2, we adopt the ablation setup of training for 100 epochs, compared to the full training of 300 epochs. The experiments are based on ViT-S/16 (Dosovitskiy et al., 2021) in DeiT (Touvron et al., 2021) with batch size 1024, which are also used in other ablation experiments in our project. The whole training process involves 20 epochs of warm-up and the remaining 80 epochs adopt the cosine annealing learning rate schedule (Loshchilov & Hutter, 2016). Each experiment takes around 2 days on 4$\times$×A100 GPUs.

### v2

@@ B.5 Breakdown Metrics for 3D VQA
We additionally provide the full metrics on SQA3D in Table E, supplementing Table 5(b). As shown in the table, adding a frozen transformer improves the performance on the main metric and most of the analytical metrics.
@@
SQA3D-LLaMA 48.09 89.03 34.27 67.05 48.17 68.34 43.87 45.64
Table E: Performance of SQA3D and adding language transformers on 3D question answering (QA). Adding LLaMA achieves the best performance. EM@1 and EM@10 means Top-1 and Top-10 Exact Match (Accuracy) metric. “What,” “Is,” “How,” “Can,” “Which,” and “Others” are detailed breakdown of question types reported in EM@1.
C Implementation Details
@@ C.2 Point Cloud Classification
In this section, we describe the implementation details of the point cloud classification method presented in Sec. 4.2. For optimization, we use the AdamW (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) optimizer and a cosine annealing learning rate schedule (Loshchilov & Hutter, 2016). To map the dimension between PointBERT and LLaMa transformer tokens, we add two linear layers with a learning rate of 5e-5. The PointBERT backbone has a learning rate of 5e-4, consistent with the original setting in Yu et al. (2021). We fine-tune our model for 300 epochs on both the ScanObjectNN (Uy et al., 2019) and ModelNet40 (Goyal et al., 2021) datasets. The training takes around 6-10 hours on 4$\times$×A100 GPUs
@@ C.3 Action Recognition
We investigate action recognition in Sec. 4.3 and provide more details of implementation here. Our setup strictly follows VideoMAE (Tong et al., 2022), where a ViT model is (1) pre-trained by masked auto-encoding (He et al., 2022); then (2) fine-tuned for additional epochs. As stated in Sec. 4.3, we directly begin from the second step and inherit the parameters of self-attention blocks in ViT from pre-trained VideoMAE models. During the training process, VideoMAE trains ViT-S for 40 epochs (5 epochs of warm-up) and ViT-B for 30 epochs (5 epochs of warm-up), where we adopt the same length of training. The optimizer is AdamW (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with a cosine annealing learning rate schedule (Loshchilov & Hutter, 2016).
In Sec. 4.3 and Table 3, our ViT-S and ViT-B performance is lower than the reported numbers in VideoMAE. This is because VideoMAE uses 32$\sim$64 GPUs during the fine-tuning stage and supports a much larger batch size compared with ours, though we scale the learning rate according to the batch size as Goyal et al. (2017a). Concretely, VideoMAE adopts the batch size of 384 video clips, while our computational resource only supports a batch size of 24 clips and 12 clips for ViT-S and ViT-B, respectively. However, we control the setup between ViT and ViT-LLaMA for a fair comparison. Finally, the ViT-S-LLaMA and ViT-B-LLaMA experiments take around 3-4 days on 4$\times$×A100 GPUs.
@@ C.4 Motion Forecasting
Table F: Our implementation of VectorNet is better than their paper. Please note that this table uses the same single-modal setting (k=1) as VectorNet for a fair comparison.
We evaluate the effects of frozen LLM transformers in Sec. 4.4 with motion forecasting. For clarity, we intuitively demonstrate its problem setting and modular architecture in Fig. G, where VectorNet (Gao et al., 2020) or mmTransformer (Liu et al., 2021) encodes each lane or agent trajectory into a token embedding, and then our LLM blocks process these tokens and feed them into the regression decoder.
Since VectorNet and mmTransformer have not released their training code, we reproduce their results on our own and achieve better or similar results as reported in their papers. As in Table F and Table G, the baselines used in our paper (Table 4) have comparable or even better performance compared with their original performance in the papers, which is critical for a fair and meaningful comparison.
@@
Table G: Our implementation of mmTransformer is comparable to the performance in their paper, with large advantages on the main metric of FDE.
During the training time, we separately train VectorNet or mmTransformer. VectorNet is a relatively simple architecture, so its training lasts 60 epochs, with a cosine annealing learning rate schedule (Loshchilov & Hutter, 2016). We use the AdamW optimizer (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with a learning rate of 5e-4 and a batch size equal to 32 samples. For mmTransformer, we train it with the same learning rate, batch size, and optimizer as VectorNet. The training lasts 32 epochs, where we drop the learning rate by 1/4 on epochs 20, 24, and 28. The training time for both models is around 2 days on one A100 GPU.
Figure G: Illustration of the typical motion forecasting design and our implementation. Motion forecasting models the trajectories of agents and lanes as polylines. Exiting models (Gao et al., 2020; Liu et al., 2021) use MLPs or transformers to convert the lanes and agent trajectories into token embeddings, and then employ a decoder to regress the future trajectories. In our design, we treat either VectorNet (Gao et al., 2020) or mmTransformer (Liu et al., 2021) as a general encoder, and then insert the frozen LLM blocks to process their embeddings.
@@
This section provides more details on implementation and designs to supplement our discussion on 2D vision-language models in Sec. 4.5. Our experiments adopt the widely-used METER (Dou et al., 2022) as the baseline and incorporate pre-trained LLM transformers after its vision-language fusion module. In Fig. H, we intuitively illustrate the modular design of METER and the specific place to insert our frozen LLM transformer and linear layers.
Conventionally, METER follows a two-stage training strategy: (1) pre-training the whole vision-language model (VLM) on a large combination of vision-language datasets, including COCO (Lin et al., 2014), Conceptual Captions (Sharma et al., 2018), SBU Captions (Ordonez et al., 2011), and Visual Genome (Krishna et al., 2017); (2) fine-tuning on downstream tasks like visual question answering (VQA) or image-text retrieval. However, the first step of pre-training is computationally extensive, so we adopt the setup in Shi et al. (2023) by skipping the pre-training step and directly training on the target task. Specifically, we initialize the image encoder from CLIP-B/32 (Radford et al., 2021) and text encoder from RoBERTa (Liu et al., 2019), and then fine-tune all the modules jointly expect for the LLM transformer $\mathbf{F}_{LM}$. Because of the initialization from CLIP and RoBERTa, our model is capable of predicting reasonably.
During the training stage, we strictly follow the same hyper-parameters and configurations on VQAv2 (Goyal et al., 2017c) and Flickr30k (Plummer et al., 2015) provided by METER. The most critical detail is that METER assigns different learning rates for each module. For example, the cross-modal fusion module and the decoder have larger learning rates compared with the pre-trained image and text encoders. Similarly, our experiments set the learning rates of linear layers ($\mathbf{F}_{L}^{1}$ and $\mathbf{F}_{L}^{2}$) 10$\times$ the learning rate of the image encoder, because they are randomly initialized. Finally, each training on VQAv2 and Flickr30k lasts for 10 epochs and around 1 day on 4$\times$×A100 GPUs.
Figure H: Illustration of the METER (Dou et al., 2022) framework and how to insert the frozen language transformer (pink $\mathbf{F}_{LM}$) and linear layers (green $\mathbf{F}_{L}^{1}$ and $\mathbf{F}_{L}^{2}$) to process the visual tokens after vision-language fusion.
@@
This section provides more details of the dataset and training configurations of the 3D vision-language task. We conduct our experiments on the SQA3D dataset (Ma et al., 2023), which contains $33.4k$ questions in $650$ unique ScanNet scenes. In addition to question answering (QA), the benchmark also requires the model to understand its situation (position, orientation, etc.) in the 3D scene as described by text. Hence it is called situated question answering (SQA). We use a batch size of 32 during our model training, and AdamW as our optimizer. The hidden size for each embedding token is 768. We train all parameters from scratch for 30 epochs, and decrease the learning rate by 10 times at 10, 15, and 20-th epoch. The model is trained on a single A100 GPU.
@@ C.7 Depths of LLM Layers
When varying the depth of transformer blocks in Sec. 5.2 and Fig. 2, we adopt the ablation setup of training for 100 epochs, compared with the full training of 300 epochs. The experiments are based on ViT-S/16 (Dosovitskiy et al., 2021) in DeiT (Touvron et al., 2021) with a batch size of 1,024, which is also used in other ablation experiments. The whole training process involves 20 epochs of warm-up and the remaining 80 epochs adopt a cosine annealing learning rate schedule (Loshchilov & Hutter, 2016). Each experiment takes around 2 days on 4$\times$×A100 GPUs.